OpenStudy (anonymous):

6√x^9.can u please tell me the steps in differentiating this question

OpenStudy (cwrw238):

first write it as f(x) = 6 x ^(9/2) now use the power rule for differentiating: if f(x) = ax^n then f'(x) = anx^(n-1) so for f(x) = 6 x ^(9/2) f'(x) = 6*(9/2) x ^ (9/2 - 1) = 27 x ^(7/2) which we can write in the form 27√x^7

OpenStudy (cwrw238):

hope this hlps

OpenStudy (cwrw238):

so does polly and bonny (on the pic)

OpenStudy (anonymous):

is this really \[6\sqrt{x^9}\] or \[\sqrt[6]{x^9}\]

OpenStudy (cwrw238):

oh - good point!!

OpenStudy (anonymous):

its the first option

OpenStudy (anonymous):

cwrw238 thank so so much..i had the answer just needed the steps

OpenStudy (cwrw238):

yw